What temperature in 0C must 7.5 L of gas at 0 0C be heated to expand its volume to 5.25 liters?


1
Expert's answer
2021-10-10T10:17:54-0400

V1/T1= V2/T2

7.5/X= 5.25/273

7.5/X= 0.019

0.019x = 7.5

× = 394

= 394.7-273

= 121.7°C